# Calogero–Sutherland hyperbolic system and Heckman–Opdam $\mathfrak{gl}_{n}$ hypergeometric function

### v1

@@ 1 Introduction
$\Psi_{{\bm{\lambda}}_{n}}(\bm{x}_{n})=\prod_{k=1}^{n-1}\frac{(2\pi\Gamma(g))^{k}}{k!}\,\int_{\mathbb{R}^{n(n-1)/2}}\exp\biggl{(}2\pi\imath\sum_{k=1}^{n}\lambda_{k}\biggl{(}\sum_{i=1}^{k}y_{ki}-\sum_{i=1}^{k-1}y_{k-1,i}\biggr{)}\biggr{)}\\[6.0pt] \times\prod_{k=1}^{n-1}\frac{\prod_{1\leq i<j\leq k}\,[4\operatorname{sh}^{2}\pi(y_{ki}-y_{kj})]^{g}}{\prod_{i=1}^{k+1}\prod_{j=1}^{k}\,[2\operatorname{ch}\pi(y_{k+1,i}-y_{kj})]^{g}}\,\prod_{k=1}^{n-1}\prod_{i=1}^{k}dy_{ki},$ (1.2)
@@
$\hat{\Psi}_{{\bm{\lambda}}_{n}}(\bm{x}_{n})=\prod_{k=1}^{n-1}\frac{(2\pi\Gamma(g))^{-k}}{k!}\,\int_{\mathbb{R}^{n(n-1)/2}}\exp\biggl{(}2\pi\imath\sum_{k=1}^{n}x_{k}\biggl{(}\sum_{i=1}^{k}\gamma_{ki}-\sum_{i=1}^{k-1}\gamma_{k-1,i}\biggr{)}\biggr{)}\\[6.0pt] \times\prod_{k=1}^{n-1}\frac{\prod_{i=1}^{k+1}\prod_{j=1}^{k}\,\Gamma\bigl{(}\imath\gamma_{k+1,i}-\imath\gamma_{kj}+\frac{g}{2}\bigr{)}\,\Gamma\bigl{(}\imath\gamma_{kj}-\imath\gamma_{k+1,i}+\frac{g}{2}\bigr{)}}{\prod_{1\leq i\not=j\leq k}\,\Gamma(\imath\gamma_{ki}-\imath\gamma_{kj})\Gamma(\imath\gamma_{kj}-\imath\gamma_{ki}+g)}\,\prod_{k=1}^{n-1}\prod_{i=1}^{k}d\gamma_{ki},$ (1.3)
where we identify $\gamma_{ni}\equiv\lambda_{i}$, was found in [KK1]. Besides, there is a well developed theory of Heckman–Opdam hypergeometric functions [HO, O], which analyses wave functions through their series representation. However, the precise correspondence between different parts of the story was not rigorously established, except for the case of two particles [BDKK6], and the goal of this paper is to fill this gap.
@@
The hyperbolic Ruijsenaars model is self-dual in a sense that corresponding wave functions diagonalize the same difference operators with respect to coordinates and spectral parameters [BDKK2]. In the case of Calogero–Sutherland system the dual model on the space of spectral parameters is governed by rational Ruijsenaars–Macdonald operators, which we consider in Section 4. The corresponding limiting procedure from hyperbolic model is different. Again, we have a pointwise limit of kernel and measure functions, but this is not a uniform limit, which can be used in integrals. Therefore here we develop the theory of Baxter operators and their properties independently. In particular, we give a direct proof of commutativity of dual Baxter operators, which is a simplified version of the hyperbolic derivation from [BDKK1]. Besides, using the Mellin–Barnes integral formula for the wave function (1.3) we describe the spectrum of dual Baxter operators.
@@
in Weyl chamber of spatial parameters $x_{1}\gg x_{2}\gg\ldots\gg x_{n}$. Comparing it with the known asymptotics of $\mathfrak{gl}_{n}$ Heckman–Opdam hypergeometric function $F_{{\bm{\lambda}}_{n}}(\bm{x}_{n})$, normalized as $F_{{\bm{\lambda}}_{n}}(0,...,0)=1$, in Section 6 we show, using the uniqueness property of Heckman–Opdam hypergeometric function [HO, O], that
@@ 2 Ruijsenaars hyperbolic system
$T^{a}_{x_{i}}=e^{a\partial_{x_{i}}},\qquad\left(T^{a}_{x_{i}}\,f\right)(x_{1},\ldots,x_{i},\ldots,x_{n})=f(x_{1},\ldots,x_{i}+a,\ldots,x_{n})$ (2.8)
@@
$Q^{R}(\bm{x}_{n},\bm{y}_{n};\lambda)=e^{{2\pi\imath}\lambda(\underline{\bm{x}}_{n}-\underline{\bm{y}}_{n})}K^{R}(\bm{x}_{n},\bm{y}_{n})\mu^{R}(\bm{y}_{n}),\qquad x_{j},y_{j}\in\mathbb{R},$ (2.20)
@@
$\Lambda^{R}(\bm{x}_{n},\bm{y}_{n-1};\lambda)=e^{{2\pi\imath}\lambda(\underline{\bm{x}}_{n}-\underline{\bm{y}}_{n-1})}K^{R}(\bm{x}_{n},\bm{y}_{n-1})\mu^{R}(\bm{y}_{n-1}),\qquad x_{j},y_{j}\in\mathbb{R},$ (2.25)
@@
is a joint eigenfunction of Macdonald operators.
@@
$\Lambda^{R}_{n}(\lambda)\Lambda^{R}_{n-1}(\rho)=\Lambda^{R}_{n-1}(\rho)\Lambda^{R}_{n}(\lambda),$ (2.28)
@@ 3 Calogero–Sutherland hyperbolic system
The proof of (3.50) copies the proof of analogous identity for the hyperbolic Calogero–Sutherland system, see e.g. [BDKK1, Section 2]. It is a direct consequence of kernel function identity [HR2, eq. (35)]:
@@ 3.1 Calogero–Sutherland limit
$\biggl{|}e^{{2\pi\imath}\lambda\underline{\bm{y}}_{n}}\prod_{a=1}^{2n}\prod_{i=1}^{n}K^{R}(z_{a}-y_{i})\mu^{R}(\bm{y}_{n})\biggr{|}\leq C(\bm{z}_{2n};g)\,\exp\biggl{(}2\pi(|\operatorname{Im}\lambda|-g)\sum_{j=1}^{n}|y_{j}|\biggr{)}.$ (3.67)
@@
    asymptotics
    $\mu^{R}(x)\sim e^{2\pi\frac{\mathrm{g}}{\omega_{1}\omega_{2}}|x|},\qquad K^{R}(x)\sim e^{-\pi\frac{\mathrm{g}}{\omega_{1}\omega_{2}}|x|},\qquad x\rightarrow\pm\infty;$ (3.71)
@@
    bounds
    $|\mu^{R}(x)|<C_{1}e^{2\pi\nu_{g}|x|},\qquad|K^{R}(x)|<C_{2}e^{-\pi\nu_{g}|x|},\qquad x\in\mathbb{R},$ (3.72)
@@
    $\int_{\mathbb{R}}dy\;e^{2\pi\imath\lambda y}K^{R}(y)=\sqrt{\omega_{1}\omega_{2}}\,S(\mathrm{g})\,\hat{K}^{R}(\lambda).$ (3.74)
@@
    $\int_{\mathbb{R}}dy\;e^{{2\pi\imath}{}\lambda y}K(y)=\frac{\Gamma\left(\imath\lambda+\frac{g}{2}\right)\Gamma\left(-\imath\lambda+\frac{g}{2}\right)}{2\pi\Gamma(g)}=\frac{1}{2\pi\Gamma(g)}\hat{K}(\lambda).$ (3.77)
@@ 4.1 Statements
First, we introduce rational analogs of weight, measure and kernel functions, see the limiting formulas (A.59). Set
@@
$\hat{Q}({\bm{\lambda}}_{n},\bm{\gamma}_{n};x)=e^{{2\pi\imath}x(\underline{\bm{\lambda}}_{n}-\underline{\bm{\gamma}}_{n})}\hat{K}({\bm{\lambda}}_{n},\bm{\gamma}_{n})\hat{\mu}(\bm{\gamma}_{n}),\qquad\lambda_{j},\gamma_{j}\in\mathbb{R}.$ (4.84)
@@
$\hat{\Lambda}({\bm{\lambda}}_{n},\bm{\gamma}_{n-1};x)=e^{{2\pi\imath}x(\underline{\bm{\lambda}}_{n}-\underline{\bm{\gamma}}_{n-1})}\hat{K}({\bm{\lambda}}_{n},\bm{\gamma}_{n-1})\hat{\mu}(\bm{\gamma}_{n-1}),\qquad\lambda_{j},\gamma_{j}\in\mathbb{R}.$ (4.88)
@@ 4.2 Proofs
$\begin{split}&\sum_{I\subset[n],\leavevmode\nobreak\ |I|=r}\prod_{i\in I}\left(\prod_{j\notin I}\frac{z_{i}-z_{j}-\alpha}{z_{i}-z_{j}}\prod_{a=1}^{n}\frac{z_{i}-y_{a}+\alpha}{z_{i}-y_{a}}\right)\\ &=\sum_{A\subset[n],\leavevmode\nobreak\ |A|=r}\prod_{a\in I}\left(\prod_{b\notin I}\frac{y_{a}-y_{b}+\alpha}{y_{a}-y_{b}}\prod_{i=1}^{n}\frac{y_{a}-z_{i}-\alpha}{y_{a}-z_{i}}\right)\end{split}$ (4.95)
@@
3. To prove Theorem 4.3 we follow the arguments of [BDKK2]. For that we first establish necessary asymptotics for dual kernel and measure functions by using well known asymptotics of gamma function
$\log\Gamma(x)=\frac{1}{2}\log 2\pi+\left(x-\frac{1}{2}\right)\log x-x+O\left(\frac{1}{x}\right),\qquad x\to\infty,\ |\arg x|\leq\pi-\delta.$
@@
$F_{1}(\bm{\gamma}_{n})=\prod_{i,j=1}^{n}\left((\lambda_{i}-\gamma_{j})^{2}+1\right)^{\frac{g-1}{2}}\prod_{1\leq i<j\leq n}\left((\gamma_{i}-\gamma_{j})^{2}+1\right)^{1-g}\prod_{j=1}^{n}\prod_{k=1}^{n-1}\left((\nu_{k}-\gamma_{j})^{2}+1\right)^{\frac{g-1}{2}},$
$\begin{split}F_{2}(\bm{\gamma}_{n},\nu_{n})=&\prod\limits_{j=1}^{n}\left(\frac{(\gamma_{j}-\nu_{n})^{2}+1}{\nu_{n}^{2}}\right)^{\frac{g-1}{2}},\qquad F_{3}(\nu_{n})=\prod\limits_{j=1}^{n-1}\left(\frac{(\nu_{j}-\nu_{n})^{2}+1}{\nu_{n}^{2}}\right)^{1-g}.\end{split}$ (4.111)
@@
Again, to perform the limit we use dominated convergence theorem. The exponential bound of the integrand coincides with that of [BDKK2] after proper change of parameters and can be chosen as in (4.109). The power part can be presented as a product

### v3

@@ 1 Introduction
$\Psi_{{\bm{\lambda}}_{n}}(\bm{x}_{n})=\prod_{k=1}^{n-1}\frac{(2\pi\Gamma(g))^{k}}{k!}\!\!\!\!\!\!\int\limits_{\mathbb{R}^{n(n-1)/2}}\!\!\!\exp\biggl(2\pi\imath\sum_{k=1}^{n}\lambda_{k}\biggl(\sum_{i=1}^{k}y_{ki}-\sum_{i=1}^{k-1}y_{k-1,i}\biggr)\biggr)\\[6.0pt] \times\prod_{k=1}^{n-1}\frac{\prod_{1\leq i<j\leq k}\,[4\operatorname{sh}^{2}\pi(y_{ki}-y_{kj})]^{g}}{\prod_{i=1}^{k+1}\prod_{j=1}^{k}\,[2\operatorname{ch}\pi(y_{k+1,i}-y_{kj})]^{g}}\,\prod_{k=1}^{n-1}\prod_{i=1}^{k}dy_{ki},$ (1.2)
@@
$\hat{\Psi}_{{\bm{\lambda}}_{n}}(\bm{x}_{n})=\prod_{k=1}^{n-1}\frac{(2\pi\Gamma(g))^{-k}}{k!}\!\!\!\!\!\!\int\limits_{\mathbb{R}^{n(n-1)/2}}\!\!\!\!\!\exp\biggl(2\pi\imath\sum_{k=1}^{n}x_{k}\biggl(\sum_{i=1}^{k}\gamma_{ki}-\sum_{i=1}^{k-1}\gamma_{k-1,i}\biggr)\biggr)\\[6.0pt] \times\prod_{k=1}^{n-1}\frac{\prod\limits_{i=1}^{k+1}\prod\limits_{j=1}^{k}\,\Gamma\bigl(\imath\gamma_{k+1,i}-\imath\gamma_{kj}+\frac{g}{2}\bigr)\,\Gamma\bigl(\imath\gamma_{kj}-\imath\gamma_{k+1,i}+\frac{g}{2}\bigr)}{\prod_{1\leq i\not=j\leq k}\,\Gamma(\imath\gamma_{ki}-\imath\gamma_{kj})\Gamma(\imath\gamma_{kj}-\imath\gamma_{ki}+g)}\,\prod_{k=1}^{n-1}\prod_{i=1}^{k}d\gamma_{ki},$ (1.3)
where we identify $\gamma_{ni}\equiv\lambda_{i}$, was found in [KK1]. Besides, there is a well developed theory of Heckman–Opdam hypergeometric functions [HO, O1], which analyses wave functions through their series representation. However, the precise correspondence between different parts of the story was not rigorously established, except for the case of two particles [BDKK6], and the goal of this paper is to fill this gap.
@@
The hyperbolic Ruijsenaars model is self-dual in a sense that corresponding wave functions diagonalize the same difference operators with respect to coordinates and spectral parameters [BDKK2]. In the case of hyperbolic Calogero–Sutherland system the dual model on the space of spectral parameters is governed by rational Ruijsenaars–Macdonald operators, which we consider in Section 4. The corresponding limiting procedure from hyperbolic model is different. Again, we have a pointwise limit of kernel and measure functions, but this is not a uniform limit, which can be used in integrals. Therefore here we develop the theory of Baxter operators and their properties independently. In particular, we give a direct proof of commutativity of dual Baxter operators, which is a simplified version of the hyperbolic derivation from [BDKK1]. Besides, using the Mellin–Barnes integral formula for the wave function (1.3) we describe the spectrum of dual Baxter operators.
@@
in Weyl chamber of spatial parameters $x_{1}\gg x_{2}\gg\ldots\gg x_{n}$. Comparing it with the known asymptotics of $\mathfrak{gl}_{n}$ Heckman–Opdam hypergeometric function $F_{{\bm{\lambda}}_{n}}(\bm{x}_{n})$, normalized as $F_{{\bm{\lambda}}_{n}}(0,...,0)=1$, in Section 6 we show, using the uniqueness property of Heckman–Opdam hypergeometric function [HO, O1], that
@@ 2 Ruijsenaars hyperbolic system
$T^{a}_{x_{i}}=e^{a\partial_{x_{i}}},\qquad\left(T^{a}_{x_{i}}\,f\right)(x_{1},\ldots,x_{n})=f(x_{1},\ldots,x_{i}+a,\ldots,x_{n})$ (2.2)
@@
$Q^{R}(\bm{x}_{n},\bm{y}_{n};\lambda)=e^{{2\pi\imath}\lambda(\underline{\bm{x}}_{n}-\underline{\bm{y}}_{n})}K^{R}(\bm{x}_{n},\bm{y}_{n})\mu^{R}(\bm{y}_{n}),$ (2.14)
@@
$\Lambda^{R}(\bm{x}_{n},\bm{y}_{n-1};\lambda)=e^{{2\pi\imath}\lambda(\underline{\bm{x}}_{n}-\underline{\bm{y}}_{n-1})}K^{R}(\bm{x}_{n},\bm{y}_{n-1})\mu^{R}(\bm{y}_{n-1}),$ (2.19)
@@
is a joint eigenfunction of Macdonald operators:
@@
Let us remark that due to symmetry of the double sine function $S_{2}(z|\omega_{1},\omega_{2})=S_{2}(z|\omega_{2},\omega_{1})$, the function $\Psi^{R}_{{\bm{\lambda}}_{n}}(\bm{x}_{n})$ is symmetric with respect to periods. Hence, it also diagonalizes the Macdonald operators with interchanged periods $\omega_{1}\leftrightarrows\omega_{2}$.
@@
$\Lambda^{R}_{n}(\lambda)\Lambda^{R}_{n-1}(\rho)=\Lambda^{R}_{n}(\rho)\Lambda^{R}_{n-1}(\lambda),$ (2.22)
@@
Due to this property and also symmetry with respect to the periods $\omega_{1}\leftrightarrows\omega_{2}$, wave functions not only diagonalize the operators (2.1), but also
$\hat{M}^{R}_{r}:=\sum_{\begin{subarray}{c}I\subset[n]\\ |I|=r\end{subarray}}\prod_{\begin{subarray}{c}i\in I\\ j\notin I\end{subarray}}\frac{\operatorname{sh}\pi\omega_{2}\bigl(\lambda_{i}-\lambda_{j}-\imath\frac{\omega_{1}+\omega_{2}-\mathrm{g}}{\omega_{1}\omega_{2}}\bigr)}{\operatorname{sh}\pi\omega_{2}(\lambda_{i}-\lambda_{j})}\cdot T^{-\imath/\omega_{1}}_{I,\lambda},$ (2.33)
that is
$\hat{M}^{R}_{r}\,\Psi^{R}_{{\bm{\lambda}}_{n}}(\bm{x}_{n})=e_{r}\bigl(e^{2\pi x_{1}/\omega_{1}},\ldots,e^{2\pi x_{n}/\omega_{1}}\bigr)\,\Psi^{R}_{{\bm{\lambda}}_{n}}(\bm{x}_{n}).$ (2.34)
@@ 3 Calogero–Sutherland hyperbolic system
The proof of (3.12) copies the proof of analogous identity for the hyperbolic Ruijsenaars system, see e.g. [BDKK1, Section 2]. It is a direct consequence of kernel function identity [HR2, eq. (35)]:
@@ 3.1 Calogero–Sutherland limit
$\biggl|e^{{2\pi\imath}\lambda\underline{\bm{y}}_{n}}$ $\prod_{a=1}^{2n}\prod_{i=1}^{n}K^{R}(z_{a}-y_{i})\mu^{R}(\bm{y}_{n})\biggr|\leq C(\bm{z}_{2n};g)\,\exp\biggl(2\pi(|\operatorname{Im}\lambda|-g)\sum_{j=1}^{n}|y_{j}|\biggr).$
@@
    asymptotics for $x\rightarrow\pm\infty$
    $\mu^{R}(x)\sim e^{2\pi\frac{\mathrm{g}}{\omega_{1}\omega_{2}}|x|},\qquad K^{R}(x)\sim e^{-\pi\frac{\mathrm{g}}{\omega_{1}\omega_{2}}|x|};$ (3.32)
@@
    bounds for $x\in\mathbb{R}$
    $|\mu^{R}(x)|<C_{1}e^{2\pi\nu_{g}|x|},\qquad|K^{R}(x)|<C_{2}e^{-\pi\nu_{g}|x|},$ (3.33)
@@
    $\int_{\mathbb{R}}dy\;e^{2\pi\imath\lambda y}K^{R}(y)=\sqrt{\omega_{1}\omega_{2}}\,S(\mathrm{g})\,\hat{K}^{R}(\lambda).$
@@
    $\int_{\mathbb{R}}dy\;e^{{2\pi\imath}{}\lambda y}K(y)=\frac{\Gamma\left(\imath\lambda+\frac{g}{2}\right)\Gamma\left(-\imath\lambda+\frac{g}{2}\right)}{2\pi\Gamma(g)}=\frac{1}{2\pi\Gamma(g)}\hat{K}(\lambda).$
@@ 4.1 Statements
Remark 4.1.
Recall that the hyperbolic wave functions diagonalize the operators (2.33)
$\hat{M}^{R}_{r}=\sum_{\begin{subarray}{c}I\subset[n]\\ |I|=r\end{subarray}}\prod_{\begin{subarray}{c}i\in I\\ j\notin I\end{subarray}}\frac{\operatorname{sh}\pi\omega_{2}\bigl(\lambda_{i}-\lambda_{j}-\imath\frac{\omega_{1}+\omega_{2}-\mathrm{g}}{\omega_{1}\omega_{2}}\bigr)}{\operatorname{sh}\pi\omega_{2}(\lambda_{i}-\lambda_{j})}\cdot T^{-\imath/\omega_{1}}_{I,\lambda}.$ (4.3)
The latter tend to the rational Hamiltonians (4.1) in the nonrelativistic limit from Section 3.1: taking $\omega_{1}=1$, $\mathrm{g}=g\omega_{2}$ and $\omega_{2}\to 0$. This is why we consider the rational Ruijsenaars system with reflected coupling constant 1g1-g. Besides, let us note that the rational Hamiltonians with constants $g$ and 1g1-g are related by similarity transformation, see (6.14).
Now we introduce rational analogs of weight, measure and kernel functions, see the limiting formulas (A.14). Set
@@
$\hat{Q}({\bm{\lambda}}_{n},\bm{\gamma}_{n};x)=e^{{2\pi\imath}x(\underline{\bm{\lambda}}_{n}-\underline{\bm{\gamma}}_{n})}\hat{K}({\bm{\lambda}}_{n},\bm{\gamma}_{n})\hat{\mu}(\bm{\gamma}_{n}).$ (4.8)
@@
$\hat{\Lambda}({\bm{\lambda}}_{n},\bm{\gamma}_{n-1};x)=e^{{2\pi\imath}x(\underline{\bm{\lambda}}_{n}-\underline{\bm{\gamma}}_{n-1})}\hat{K}({\bm{\lambda}}_{n},\bm{\gamma}_{n-1})\hat{\mu}(\bm{\gamma}_{n-1}).$ (4.12)
@@ 4.2 Proofs
$\begin{split}&\sum_{I\subset[n],\penalty 10000\ |I|=r}\prod_{i\in I}\left(\prod_{j\notin I}\frac{z_{i}-z_{j}-\alpha}{z_{i}-z_{j}}\prod_{a=1}^{n}\frac{z_{i}-y_{a}+\alpha}{z_{i}-y_{a}}\right)\\ &=\sum_{A\subset[n],\penalty 10000\ |A|=r}\prod_{a\in A}\left(\prod_{b\notin A}\frac{y_{a}-y_{b}+\alpha}{y_{a}-y_{b}}\prod_{i=1}^{n}\frac{y_{a}-z_{i}-\alpha}{y_{a}-z_{i}}\right)\end{split}$ (4.19)
@@
3. To prove Theorem 4.3 we follow the arguments of [BDKK2]. For that we first establish necessary asymptotics for dual kernel and measure functions by using well known asymptotics of gamma function when $x\to\infty,\ |\arg x|\leq\pi-\delta$:
$\log\Gamma(x)=\frac{1}{2}\log 2\pi+\left(x-\frac{1}{2}\right)\log x-x+O\left(\frac{1}{x}\right).$
@@
$\begin{split}F_{1}(\bm{\gamma}_{n})=\prod_{i,j=1}^{n}\left((\lambda_{i}-\gamma_{j})^{2}+1\right)^{\frac{g-1}{2}}\prod_{1\leq i<j\leq n}\left((\gamma_{i}-\gamma_{j})^{2}+1\right)^{1-g}\;\prod_{j=1}^{n}\prod_{k=1}^{n-1}\left((\nu_{k}-\gamma_{j})^{2}+1\right)^{\frac{g-1}{2}},\end{split}$
$\begin{split}F_{2}(\bm{\gamma}_{n},\nu_{n})=&\prod\limits_{j=1}^{n}\left(\frac{(\gamma_{j}-\nu_{n})^{2}+1}{\nu_{n}^{2}}\right)^{\frac{g-1}{2}},\qquad F_{3}(\nu_{n})=\prod\limits_{j=1}^{n-1}\left(\frac{(\nu_{j}-\nu_{n})^{2}+1}{\nu_{n}^{2}}\right)^{1-g}.\end{split}$
@@
Again, to perform the limit we use dominated convergence theorem. The exponential bound of the integrand coincides with that given in [BDKK2] after proper change of parameters and can be chosen as in (4.33). The power part can be presented as a product